OpenStudy (samirahdanyel):

Which statements are true regarding the area of circles and sectors? Check all that apply. The area of a circle depends on the length of the radius. The area of a sector depends on the ratio of the central angle to the entire circle. The area of a sector depends on pi. The area of the entire circle can be used to find the area of a sector. The area of a sector can be used to find the area of a circle.

OpenStudy (photon336):

\[A_{circle} = \pi*r^{2}\]

OpenStudy (samirahdanyel):

so its the first one?

OpenStudy (photon336):

based on that formula does the area of the circle depend on the length of the radius?

OpenStudy (samirahdanyel):

noo

OpenStudy (samirahdanyel):

i dont know

OpenStudy (photon336):

look at the formula

OpenStudy (photon336):

it will make sense, let me show you say if the radius is 4 and we double it to 8

OpenStudy (samirahdanyel):

okay

OpenStudy (photon336):

\[\pi*(4)^{2} = 16\pi\] \[\pi*(8)^{2} = 64\pi\] take a look at the two radii 4 and 8 look what happens to the area? can you tell me how radius affects area?

OpenStudy (samirahdanyel):

the area increases

OpenStudy (photon336):

yep and why?

OpenStudy (samirahdanyel):

because the radi increases

OpenStudy (photon336):

exactly so what can we say about A

OpenStudy (photon336):

for an area of a sector where theta is the angle made by the sector \[\frac{ \theta }{ 360 }*\pi*r^{2}\]

OpenStudy (photon336):

well the area of a sector depends on pi it also depends on r too.

OpenStudy (photon336):

pi*r^2 is the area of the entire circle so yeah it can be used to find the area of a sector, same with the area of the sector.